# Solving Linear Systems of Equations with the Quantum HHL Algorithm: A Tutorial on the Physical and Mathematical Foundations for Undergraduate Students

### v1

@@ Appendix B Computational Complexity Theory
Asymptotic analysis is an essential tool for understanding the efficiency of algorithms, as it focuses on the behavior of complexity functions when the input size is sufficiently large Papadimitriou (2003). Big-O notation, defined as $O(f(n))$, indicates that the algorithm’s complexity does not grow faster than $f(n)$ for sufficiently large inputs. In other words, the time or space required by the algorithm is at most proportional to $f(n)$ in large-input cases Alsuwaiyel (2003). On the other hand, Omega notation, represented by $\Omega(f(n))$, indicates that the algorithm’s complexity has a lower bound, ensuring that the algorithm will require at least resources proportional to $f(n)$ for sufficiently large inputs Alsuwaiyel (2003). Theta notation, or $\Theta(f(n))$, shows that the algorithm’s complexity grows exactly in the same order as $f(n)$, being neither greater nor smaller than $f(n)$ for large inputs Alsuwaiyel (2003).
These notations allow algorithms to be grouped into distinct complexity classes. For example, when the runtime or space required does not vary with input size, we say it has constant complexity, or $O(1)$. When the growth follows logarithmic, linear, linear-logarithmic, quadratic, exponential, or factorial patterns, we respectively use the notations $O(\log n)$, $O(n)$, $O(n\log n)$, $O(n^{2})$, $O(2^{n})$, and $O(n!)$ Alsuwaiyel (2003). Each of these categories reflects how the required resources increase as the problem grows, making them a fundamental part of algorithm evaluation. The relationship between these complexities is illustrated in Fig. 8.
@@
In this context, an algorithm can be considered efficient if the execution time does not scale faster than a given polynomial function of the input size Ekert and Jozsa (1996). In general, the input size can be given by the total number of bits representing it, while the execution time is measured through the number of computational steps required Nielsen and Chuang (2010); Ekert and Jozsa (1996). Thus, a complexity class can be described as a set of languages defined by a complexity measure, such as the number of computational resources used, the number of queries required, or simply the runtime, which relates a given string $\{l\}$ to a language $L$ Cormen (2012). Among them, the $P$ complexity class refers to a set of decision problems that can be solved in polynomial time, while the $NP$ class corresponds to languages that can be verified by a polynomial-time algorithm Nielsen and Chuang (2010); Papadimitriou (2003); Cormen (2012).
While class $P$ considers solving problems in polynomial time by a deterministic Turing machine, the class PSPACE encompasses problems solvable in polynomial space regardless of the time consumed. Class $NP$ refers to solving problems in polynomial time but for which no efficient solution method is known. $NP$ contains $NP$-complete problems, which are characterized as the hardest problems in this class. If any $NP$-complete problem can be solved in polynomial time, then all problems in $NP$ can be solved in polynomial time as well Arora and Barak (2009). The $BQP$ class (Bounded-Error Quantum Polynomial Time) refers to problems that can be solved by quantum computers in polynomial time, with a bounded probability of error. $BQP$ is the class addressing quantum algorithms, analogous to the classical $BPP$ class, where probabilistic algorithms with bounded error are allowed. A summary of these classes is shown in Fig. 9.
@@
In this case, quantum algorithms can efficiently solve probabilistic classical algorithms but face restrictions in solving $NP$-complete problems in polynomial time Arora and Barak (2009).

### v2

@@ Appendix B Computational Complexity Theory
Asymptotic analysis is an essential tool for understanding the efficiency of algorithms, as it focuses on the behavior of complexity functions when the input size is sufficiently large papadimitriou2003computational. Big-O notation, defined as $O(f(n))$, indicates that the algorithm’s complexity does not grow faster than $f(n)$ for sufficiently large inputs. In other words, the time or space required by the algorithm is at most proportional to $f(n)$ in large-input cases alsuwaiyel2003. On the other hand, Omega notation, represented by $\Omega(f(n))$, indicates that the algorithm’s complexity has a lower bound, ensuring that the algorithm will require at least resources proportional to $f(n)$ for sufficiently large inputs alsuwaiyel2003. Theta notation, or $\Theta(f(n))$, shows that the algorithm’s complexity grows exactly in the same order as $f(n)$, being neither greater nor smaller than $f(n)$ for large inputs alsuwaiyel2003.
These notations allow algorithms to be grouped into distinct complexity classes. For example, when the runtime or space required does not vary with input size, we say it has constant complexity, or $O(1)$. When the growth follows logarithmic, linear, linear-logarithmic, quadratic, exponential, or factorial patterns, we respectively use the notations $O(\log n)$, $O(n)$, $O(n\log n)$, $O(n^{2})$, $O(2^{n})$, and $O(n!)$ alsuwaiyel2003. Each of these categories reflects how the required resources increase as the problem grows, making them a fundamental part of algorithm evaluation. The relationship between these complexities is illustrated in Fig. 8.
@@
In this context, an algorithm can be considered efficient if the execution time does not scale faster than a given polynomial function of the input size ekert1996quantum. In general, the input size can be given by the total number of bits representing it, while the execution time is measured through the number of computational steps required nielsen2010quantum; ekert1996quantum. Thus, a complexity class can be described as a set of languages defined by a complexity measure, such as the number of computational resources used, the number of queries required, or simply the runtime, which relates a given string $\{l\}$ to a language $L$ cormen2012algoritmos. Among them, the $P$ complexity class refers to a set of decision problems that can be solved in polynomial time, while the $NP$ class corresponds to languages that can be verified by a polynomial-time algorithm nielsen2010quantum; papadimitriou2003computational; cormen2012algoritmos.
While class $P$ considers solving problems in polynomial time by a deterministic Turing machine, the class PSPACE encompasses problems solvable in polynomial space regardless of the time consumed. Class $NP$ refers to solving problems in polynomial time but for which no efficient solution method is known. $NP$ contains $NP$-complete problems, which are characterized as the hardest problems in this class. If any $NP$-complete problem can be solved in polynomial time, then all problems in $NP$ can be solved in polynomial time as well arora2009computational. The $BQP$ class (Bounded-Error Quantum Polynomial Time) refers to problems that can be solved by quantum computers in polynomial time, with a bounded probability of error. $BQP$ is the class addressing quantum algorithms, analogous to the classical $BPP$ class, where probabilistic algorithms with bounded error are allowed. A summary of these classes is shown in Fig. 9.
@@
In this case, quantum algorithms can efficiently solve probabilistic classical algorithms but face restrictions in solving $NP$-complete problems in polynomial time arora2009computational.